OpenStudy (sambam):

Please help me simplify the expression: (x^2-7x+12)/(4x^2-40x+96)*(2x^2-6x-36)/(x^2-9)

OpenStudy (zehanz):

First, you will have to factor every trinomial and also the binomial x²-9. I will do the first one for yoiu: x²-7x+12=(x+a)(x+b). Here, a + b = -7 and a*b = 12, so I take a = -3 and b = -4: -3 + -4 = -7 OK -3 * -4 = 12 OK So x²-7x+12=(x-3)(x-4). Now do all of the other ones in the same way (just factor out 4 first in 4x²-40x+96, this will make it easier). After you've done all this, you'll see lots of common factors - cancel them! There won't be much left, I promise!